The equilibrium constant, Kp, for the following reaction is 2.01 at 500 K:


PCl3(g) + Cl2(g) PCl5(g)


Calculate the equilibrium partial pressures of all species when PCl3 and Cl2, each at an intitial partial pressure of 1.47 atm, are introduced into an evacuated vessel at 500 K.


1
Expert's answer
2022-02-14T17:02:03-0500

Reaction is

PCl3 + Cl2 = PCl5 .

Initial 1.47 1.47

Equilibrium 1.47 -y 1.47 -y y


Kp = Ppcl5 / Ppcl3 × Pcl2 .

Kp = y / ( 1.47 - y ) × ( 1.47 - y ) .

2.01 = y / ( 1.47 - y )2 .


Then find y after solving above equation.

And find partial pressure y . And 1.47 - y .